# PhyMPGN: Physics-encoded Message Passing Graph Network for spatiotemporal PDE systems

### v1

@@ 1 Introduction
Many traditional numerical methods have been developed to solve PDEs (Ames, 2014; Anderson, 1995), such as finite difference methods, finite element methods, finite volume methods, spectral methods, etc. However, to achieve the targeted accuracy, the computational cost of these methods for large simulations is prohibitively high, since fine meshes and small time stepping are required. In recent years, deep learning methods have achieved great success in various domains such as image recognition (He et al., 2016), natural language processing (Vaswani et al., 2017) and information retrieval (Devlin et al., 2019). A increasing number of neural-based methods have been proposed to learn the solution of PDEs.
Physical-informed neural networks (PINNs) (Raissi et al., 2019) and its variants have shown promise performance in modeling various physical systems, such as fluid dynamics (Gu et al., 2024; Raissi et al., 2020), rarefied-gas dynamics (De Florio et al., 2021), and engineer problems (Rezaei et al., 2022; Haghighat et al., 2021; Rao et al., 2021; Eshkofti & Hosseini, 2023). However, the family of PINN methods easily encounter scalability and generalizability issues when dealing with complex systems, because of their fully connected neural network architecture and the incorporation of soft physical constraints via loss regularizer. Neural operators (Lu et al., 2021; Li et al., 2021; 2023b; Wu et al., 2023), the variants of Transformer (Geneva & Zabaras, 2022; HAN et al., 2022; Li et al., 2023a), and graph neural networks (GNN) (Battaglia et al., 2018; Seo* et al., 2020; Iakovlev et al., 2021; Pfaff et al., 2021; Brandstetter et al., 2022; Ye et al., 2024) overcome the above issues to some extent; however, they require a substantial amount of training datasets to learn the mapping between infinite function spaces and temporal evolution, due to their “black-box” data-driven learning manner. The physics-encoded learning, e.g., PeRCNN (Rao et al., 2023), has the ability to learn spatiotemporal dynamics based on limited low-resolution and noisy measurement data. However, this method is inapplicable to prediction in irregular domains. Therefore, achieving accurate predictions of spatiotemporal dynamics based on limited low-resolution training data within irregular computational domains remains a challenge.
@@ 2 Related work
Autoregressive learning methods model spatiotemporal dynamics iteratively. Generally, such methods employ a neural network to extract spatial patterns and a recurrent block to model temporal evolution, such as RNN (Hochreiter & Schmidhuber, 1997; Cho et al., 2014), Transformers (Vaswani et al., 2017), or a numerical integrator. PA-DGN (Seo* et al., 2020) combines the spatial difference layer with a recurrent GNN to learn the underlying dynamics. PhyCRNet (Ren et al., 2022) employs ConvsLSTM (Shi et al., 2015) to extract spatial features and evolve over time. LED (Vlachas et al., 2022) deploys RNNs with gating mechanisms to approximate the evolution of the coarse-scale dynamics and utilizes auto-encoders to transfer the information across fine and coarse scales. The Transformer architecture combined with different embedding techniques (Geneva & Zabaras, 2022; HAN et al., 2022) for physical systems has also been explored. Besides, there are also lots of neural models (Rao et al., 2023; Brandstetter et al., 2022; Pfaff et al., 2021; Choi et al., 2023) employing numerical integrators for temporal marching.
@@
Diffusion processes on graphs (Freidlin & Wentzell, 1993; Freidlin & Sheu, 2000) have a various applications, such as image processing (Lozes et al., 2014; Gilboa & Osher, 2009) and data analysis (Coifman et al., 2005; Belkin & Niyogi, 2003). Recently, studies exploring the connection between these processes and GNNs have been increasing.Implementing diffusion operations on graphs enhances the representational capacity of graph learning (Atwood & Towsley, 2016; Liao et al., 2019; Gasteiger et al., 2019). The notion of the diffusion PDE on graphs also inspire the understanding and design GNNs (Chamberlain et al., 2021; Thorpe et al., 2022).
@@ 3.1 Problem setup
We focus on a spatial domain $\Omega$ with sparsely observed nodes $\{\bm{x}_{0},\dots,\bm{x}_{N-1}\}$, non-uniformly placed in space (e.g., on an unstructured mesh), presenting a more challenging scenario compared with structured grids. Observations $\{\bm{U}(t_{0}),\dots,\bm{U}(t_{T-1})\}$ are collected at time points $t_{0},\dots,t_{T-1}$, where $\bm{U}(t_{i})=\{\bm{u}(\bm{x}_{0},t_{i}),\dots,\bm{u}(\bm{x}_{N-1},t_{i})\}$ denotes the physical quantities. Considering that many physical phenomena involve diffusion processes, we assume the diffusion term in the PDE is known as a priori knowledge. Our goal is to develop a graph learning model with small training datasets capable of accurately predicting various spatiotemporal dynamics on coarse unstructured meshes, handling different types of BCs, and producing the trajectory of dynamics for an arbitrarily given IC.
@@ 3.2 Model architecture
Figure 1: (a) Model with the second-order Runge-Kutta scheme. (b) NN block consists of two parts: a GNN block followed the Encode-Process-Decode framework and a Laplace block with the correction architecture. Mesh Laplace module in Laplace block denotes the discrete Laplace-Beltrami operator in geometric mathematics. The other three modules, $\text{MLP}_{\alpha},\text{MLP}_{\beta}$, and MPNNs, constitutes the lightweight learnable network for correction. (c) The MPNNs module in GNN block consists of $L$ identical MPNN layers. Residual connection and padding in latent space are applied in the first $L-1$ layers, excluding the last layer.
@@ 3.2.1 GNN block
$\begin{split}\text{edge}\ (i,j)\ \text{message:}\quad&\bm{m}^{l}_{ij}=\phi\left(\bm{h}^{l}_{i},\ \bm{h}^{l}_{j}-\bm{h}^{l}_{i},\ \bm{e}_{ij}\right),\ 0\leq l<L\\ \text{node}\ i\ \text{update:}\quad&\bm{h}^{l+1}_{i}=\gamma\bigg{(}\bm{h}^{l}_{i},\ \bigoplus\limits_{j\in\mathcal{N}_{i}}\bm{m}^{l}_{ij}\bigg{)}+\bm{h}^{l}_{i},\ 0\leq l<L-1\\ &\bm{h}^{L}_{i}=\gamma\bigg{(}\bm{h}^{L-1}_{i},\ \bigoplus\limits_{j\in\mathcal{N}_{i}}\bm{m}^{L-1}_{ij}\bigg{)}\end{split}$ (3)
@@ 3.2.2 Laplace block
Figure 2: (a) Illustration of four discrete nodes ($i$, $j$, $p$, and $q$) and five edges connecting them. The two angles opposite to the edge $(i,j)$ are $\alpha_{ij}$ and $\beta_{ij}$. (b) The blue region denotes the area of the polygon formed by connecting the circumcenters (e.g., $c_{1},\ c_{2}$, etc.) of the triangles around node $i$.
@@
$\Delta f_{i}=\frac{1}{d_{i}}\Big{(}z_{i}+\sum\limits_{j\in\mathcal{N}_{i}}w_{ij}(f_{i}-f_{j})\Big{)}$ (5)
@@ 3.3 Encoding boundary conditions
Given the governing PDE, the solution depends on both ICs and BCs. Inspired by the padding methods for BCs on regular grids in PeRCNN (Rao et al., 2023), we propose a novel padding strategy to encode different types of BCs on irregular domains. Specifically, we perform BC padding in both the physical space and the latent space.
@@
Before constructing the unstructured mesh by Delaunay triangulation, we first apply the padding strategy to the discrete nodes in the physical space (e.g., $\bm{u}^{k}$), as shown in Figure 3. We consider four type of BCs: Dirichlet, Neumann, Robin, and Periodic. All nodes on the Dirichlet boundary will be padded with the specified values. For Neumann/Robin BCs, ghost nodes are created to be symmetric with the nodes nearby the boundary in the physical space, and their padded values are based on gradient information in the normal direction. For periodic BCs, the nodes near the boundary are flipped to the other side, achieving a cyclic effect in message passing. Once the padding is completed, Delaunay triangulation is applied to construct the mesh, which serves as the graph input for the model. Apart from this, we also apply padding to the prediction (e.g., $\bm{u}^{k+1}$) of the model at each time step to ensure that it satisfies the boundary conditions before being fed into the model for next-step prediction. As for the latent space (e.g., $\bm{h}^{k}$ in the GNN block), padding is also applied after each MPNN layer except the last layer. Detail formulas of encoding different types of BCs are provided in Appendix C.
@@ 3.4 Network training
Due to GPU memory limitations, directly feeding the entire long time series ($T$ steps) into the model and performing backpropagation across all steps is impractical. To address this, we segment the time series into multiple shorter time sequences ($M$ steps, $T\gg M$). The model rolls out for $M-1$ times in each segment, but backpropagation is only applied to the first and last steps. Furthermore, we introduce a small noise to the first frame of each segment during training. Both of these techniques aim to alleviate model overfitting and mitigate error accumulation. Therefore, the loss function of each time segment is defined as
$\mathcal{L}=\text{MSE}(\bm{u}^{1},\hat{\bm{u}}^{1})+\text{MSE}\left(\bm{u}^{{M-1}},\hat{\bm{u}}^{{M-1}}\right)$ (7)
where $\bm{u}$ denotes the ground truth of state variable, $\hat{\bm{u}}$ denotes the prediction from our model, and the superscript of $\bm{u},\hat{\bm{u}}$ denotes their time steps in the segment.
@@ 4 Experiments
Here, we present comprehensive evaluation of our methods. Firstly, we validate the feasibility and efficacy of the Laplace block to compute the Laplacian on coarse non-uniform meshes. Subsequently, the entire model is trained and evaluated on various PDE datasets and compared with several baseline models. Finally, we perform ablation studies to demonstrate the positive contribution of each individual component within our model and the impact of numerical integrator. Our source code and data will be posted after peer review.
@@ 4.1 Laplace block validation
We create a synthetic dataset to validate the effectiveness of the Laplace block. First, we generate 10,000 samples of random continuous functions on a high-resolution ($129\times 129$) regular grid by cumulatively adding sine and cosine functions, satisfying periodic BCs. By employing the 5-point Laplacian operator of the finite difference method, the Laplacian of the functions at each node is computed as the ground truth. Then the dataset is down-sampled in space from the high-resolution regular grids to a set of coarse non-uniform mesh points yielding only 983 nodes. Note that 70% of the dataset is used as the training set, 20% as the validation set, and 10% as the test set. More details of the synthetic dataset are shown in Appendix D.1.
We compare the performances of four different methods: (1) Laplace block with BC padding; (2) the discrete Laplace-Beltrami operator in geometric mathematics with BC padding, referred to as Mesh Laplace; (3) the spatial difference layer for Laplacian, referred to as SDL (Seo* et al., 2020); (4) SDL with BC padding, referred to as SDL-padding. The MSE and RNE of all experiments are shown in Table 2. Mesh Laplace performs poorly compared to other learnable methods, while our Laplace block with the fewest parameters greatly outperforms other methods. SDL-padding exhibits an obvious performance improvement, indicating the effectiveness and importance of our padding strategy. A snapshot of a random function, ground truth of its Laplacian, and prediction are shown in Appendix Figure A.12.
Table 1: The MSE and RNE of four different methods to approximate Laplacian on non-uniform domain.
@@ 4.2 Modeling PDE systems
Solving PDEs serves as a cornerstone for modeling complex dynamical systems. However, in fields such as climate forecasting, reactions of new chemical matters, and social networks, the underlying governing PDEs are either completely unknown or only partially known. Hence, there is a great need for data-driven modeling and simulation. Based on the assumption, we train and evaluate our model on various physical systems, including Burgers’ equation, FitzHugh-Nagumo (FN) equation, Gray-Scott (GS) equation, and flows past a cylinder, particularly in small training data regimes. We also compare our model against several representative baseline models, including DeepONet (Lu et al., 2021), PA-DGN (Seo* et al., 2020), MGN (Pfaff et al., 2021), and MP-PDE (Brandstetter et al., 2022). The numbers of learnable parameters and training iterations for all models are kept as consistent as possible to ensure a fair comparison. All the data are generated using COMSOL with fine meshes and small time stepping, which are further spatiotemporally down-sampled to establish the training and testing datasets. More details of the datasets are discussed in the Appendix D.2.
@@
Figure 6: Diagram of different domains and hybrid BCs for the four scenarios.
@@
To validate our model’s ability to encode different types of BCs, we use Burgers’ equation as an example and consider three additional scenarios with irregular domain and hybrid BCs, resulting in a total of four scenarios, as illustrated in Figure 6: (1) Scenario A, a square domain with periodic BCs, as mentioned in the previous paragraph; (2) Scenario B, a square domain with a circular cutout, where periodic and Dirichlet BCs are applied; (3) Scenario C, a square domain with a dolphin-shaped cutout, where periodic and Dirichlet BCs are applied; and (4) Scenario D, a square domain with semicircular cutouts on both the left and right sides, where periodic and Neumann BCs are applied. Similarly, 10/10 training and testing trajectories are generated and down-sampled with approximately $N=600$ observed nodes in each domain. Figure 5 depicts the error distribution over all time steps, correlation curves, and predicted snapshots at the last time step, for the last three scenarios (the results of Scenario A are presented in Figure 4a). It is seen that our model significantly outperforms other baseline models and shows the efficacy of handling PDE systems with different BCs.
@@
In all experiments, our model outperforms the baselines with considerable margins, e.g., exceeding 50% gains (see the specific numerical results reported in Appendix Tables A.6, A.8 and A.9). The predicted trajectory snapshots of our model and ground truth are provided in Appendix Figure A.14.
Figure 8: Results of the four models: Model A, B, C, and Ours.

### v2

@@ 1 Introduction
Many traditional numerical methods have been developed to solve PDEs (Ames, 2014; Anderson, 1995), such as finite difference methods, finite element methods, finite volume methods, spectral methods, etc. However, to achieve the targeted accuracy, the computational cost of these methods for large simulations is prohibitively high, since fine meshes and small time stepping are required (typically relaxed for implicit methods). In recent years, deep learning methods have achieved great success in various domains such as image recognition (He et al., 2016), natural language processing (Vaswani et al., 2017) and information retrieval (Devlin et al., 2019). A increasing number of neural-based methods have been proposed to learn the solution of PDEs.
Physical-informed neural networks (PINNs) (Raissi et al., 2019) and its variants have shown promise performance in modeling various physical systems, e.g., fluid dynamics (Gu et al., 2024; Raissi et al., 2020), rarefied-gas dynamics (De Florio et al., 2021), and engineer problems (Rezaei et al., 2022; Haghighat et al., 2021; Rao et al., 2021; Eshkofti & Hosseini, 2023). However, the family of PINN methods easily encounter scalability and generalizability issues when dealing with complex systems, due to their fully connected neural network architecture and the incorporation of soft physical constraints via loss regularizer. Neural operators (Lu et al., 2021; Li et al., 2021; 2023b; Wu et al., 2023), the variants of Transformer (Geneva & Zabaras, 2022; HAN et al., 2022; Li et al., 2023a), and graph neural networks (GNN) (Battaglia et al., 2018; Seo et al., 2020; Iakovlev et al., 2021; Pfaff et al., 2021; Brandstetter et al., 2022; Ye et al., 2024) overcome the above issues to some extent; however, they require substantial training datasets to learn the mapping between infinite function spaces and temporal evolution, due to their “black-box” data-driven learning manner. The physics-encoded learning, e.g., PeRCNN (Rao et al., 2023), has the ability to learn spatiotemporal dynamics based on limited low-resolution and noisy measurement data. However, this method is inapplicable to irregular domains. Therefore, achieving accurate predictions of spatiotemporal dynamics based on limited low-resolution training data within irregular computational domains remains a challenge.
@@ 2 Related work
Autoregressive learning methods model spatiotemporal dynamics iteratively. Generally, such methods employ a neural network to extract spatial patterns and a recurrent block to model temporal evolution, such as RNN (Hochreiter & Schmidhuber, 1997; Cho et al., 2014), Transformers (Vaswani et al., 2017), or a numerical integrator. PA-DGN (Seo et al., 2020) combines the spatial difference layer with a recurrent GNN to learn the underlying dynamics. PhyCRNet (Ren et al., 2022) employs ConvsLSTM (Shi et al., 2015) to extract spatial features and evolve over time. LED (Vlachas et al., 2022) deploys RNNs with gating mechanisms to approximate the evolution of the coarse-scale dynamics and utilizes auto-encoders to transfer the information across fine and coarse scales. The Transformer architecture combined with different embedding techniques (Geneva & Zabaras, 2022; HAN et al., 2022) for physical systems has also been explored. Besides, there are also lots of neural models (Rao et al., 2023; Brandstetter et al., 2022; Pfaff et al., 2021; Choi et al., 2023; Mi et al., 2024) employing numerical integrators for temporal marching.
@@
Diffusion processes on graphs (Freidlin & Wentzell, 1993; Freidlin & Sheu, 2000) have a various applications, such as image processing (Lozes et al., 2014; Gilboa & Osher, 2009) and data analysis (Coifman et al., 2005; Belkin & Niyogi, 2003). Recently, studies exploring the connection between these processes and GNNs have been increasing.Implementing diffusion operations on graphs enhances the representational capacity of graph learning (Atwood & Towsley, 2016; Liao et al., 2019; Gasteiger et al., 2019). The notion of the diffusion PDE on graphs also inspires the understanding and design of GNNs (Chamberlain et al., 2021; Thorpe et al., 2022).
@@ 3.1 Problem setup
We focus on a spatial domain $\Omega$ with non-uniformly and sparsely observed nodes $\{\bm{x}_{0},\dots,\bm{x}_{N-1}\}$ (e.g., on an unstructured mesh), presenting a more challenging scenario compared with structured grids. Observations $\{\bm{U}(t_{0}),\dots,\bm{U}(t_{T-1})\}$ are collected at time points $t_{0},\dots,t_{T-1}$, where $\bm{U}(t_{i})=\{\bm{u}(\bm{x}_{0},t_{i}),\dots,\bm{u}(\bm{x}_{N-1},t_{i})\}$ denote the physical quantities. Considering that many physical phenomena involve diffusion processes, we assume the diffusion term in the PDE is known as a priori knowledge. Our goal is to develop a graph learning model with small training datasets capable of accurately predicting various spatiotemporal dynamics on coarse unstructured meshes, handling different types of BCs, and producing the trajectory of dynamics for an arbitrarily given IC.
@@ 3.2 Model architecture
Figure 1: (a) Model with the second-order Runge-Kutta scheme. (b) NN block consists of two parts: a GNN block followed the Encode-Process-Decode framework and a Laplace block with the correction architecture. Mesh Laplace in Laplace block denotes the discrete Laplace-Beltrami operator in geometric mathematics. The other three modules, $\text{MLP}_{\alpha},\text{MLP}_{\beta}$, and MPNNs, constitutes the lightweight learnable network for correction. (c) The MPNNs module in GNN block consists of $L$ identical MPNN layers. Residual connection and padding in latent space are applied in the first $L-1$ layers, excluding the last layer.
@@ 3.2.1 GNN block
$\begin{split}\text{edge}\ (i,j)\ \text{message:}\quad&\bm{m}^{l}_{ij}=\phi\left(\bm{h}^{l}_{i},\ \bm{h}^{l}_{j}-\bm{h}^{l}_{i},\ \bm{e}_{ij}\right),\ 0\leq l<L\\ \text{node}\ i\ \text{update:}\quad&\bm{h}^{l+1}_{i}=\gamma\bigg{(}\bm{h}^{l}_{i},\ \bigoplus\limits_{j\in\mathcal{N}_{i}}\bm{m}^{l}_{ij}\bigg{)}+\bm{h}^{l}_{i},\ 0\leq l<L-1\\ &\bm{h}^{L}_{i}=\gamma\bigg{(}\bm{h}^{L-1}_{i},\ \bigoplus\limits_{j\in\mathcal{N}_{i}}\bm{m}^{L-1}_{ij}\bigg{)}\end{split}\vspace{-12pt}$ (3)
@@ 3.2.2 Laplace block
Figure 2: (a) Four discrete nodes ($i$, $j$, $p$, and $q$) and five edges connecting them. The two angles opposite to the edge $(i,j)$ are $\alpha_{ij}$ and $\beta_{ij}$. (b) The blue region denotes the area of the polygon formed by connecting the circumcenters (e.g., $c_{1},\ c_{2}$, etc.) of the triangles around node $i$.
@@
$\Delta f_{i}=\frac{1}{d_{i}}\Big{(}z_{i}+\sum\limits_{j\in\mathcal{N}_{i}}w_{ij}(f_{i}-f_{j})\Big{)}\vspace{-3pt}$ (5)
@@ 3.3 Encoding boundary conditions
Given the governing PDE, the solution depends on both ICs and BCs. Inspired by the padding methods for BCs on regular grids in PeRCNN (Rao et al., 2023), we propose a novel padding strategy to encode four types of BCs on irregular domains. Specifically, we perform BC padding in both the physical and latent space. PENN (Horie & MITSUME, 2022) also constructs their model to satisfy Dirichlet and Neumman BCs by designing special neural layers and modules while we apply the padding strategy directly to the features. More details of the comparison between these methods are presented in Appendix C.
@@
Before constructing the unstructured mesh by Delaunay triangulation, we first apply the padding strategy to the discrete nodes in the physical space (e.g., $\bm{u}^{k}$), as shown in Figure 3. We consider four type of BCs: Dirichlet, Neumann, Robin, and Periodic. Nodes on the Dirichlet boundary will be directly assigned specific values. For Neumann/Robin BCs, ghost nodes are created symmetrically with respect to the nodes near the boundary (along the normal direction) in the physical space, and their padded values depend on derivatives in the normal direction. The goal is to ensure the true nodes, the boundary, and the ghost nodes satisfy the BCs in a manner similar to the central difference method. For periodic BCs, the nodes near the boundary are flipped and placed near the other corresponding boundary, achieving a cyclic effect in message passing. Once the padding is completed, Delaunay triangulation is applied to construct the mesh, which serves as the graph input for the model. Apart from this, we also apply padding to the prediction (e.g., $\bm{u}^{k+1}$) of the model at each time step to ensure that it satisfies BCs before being fed into the model for next-step prediction. As for the latent space (e.g., $\bm{h}^{k}$ in the GNN block), padding is also applied after each MPNN layer except the last layer. Details of encoding different types of BCs are provided in Appendix C.
@@ 3.4 Network training
Due to GPU memory limitations, directly feeding the entire long time series ($T$ steps, $\bm{u}^{0},\dots,\bm{u}^{T-1}$) into the model and performing backpropagation across all steps is impractical. To address this, we segment the time series into multiple shorter time sequences of $M$ steps ($\bm{u}^{s_{0}},\dots,\bm{u}^{s_{M-1}}$, $T\gg M$). With the input $\bm{u}^{s_{0}}$, the model rolls out for $M-1$ times to generate predictions ($\hat{\bm{u}}^{s_{1}},\dots,\hat{\bm{u}}^{s_{M-1}}$), but backpropagation is only applied to the first and last predictions. Furthermore, we introduce a small noise to the first frame $\bm{u}^{s_{0}}$ of each segment during training. Both techniques aim to alleviate overfitting and error accumulation. Therefore, the loss function of each time segment is defined as
$\mathcal{L}=\text{MSE}(\bm{u}^{s_{1}},{{\hat{\bm{u}}}}^{s_{1}})+\text{MSE}\left(\bm{u}^{{s_{M-1}}},{\hat{\bm{u}}}^{{s_{M-1}}}\right)$ (7)
where $\bm{u}$ denotes the ground truth of state variable, $\bm{\hat{u}}$ denotes the prediction from our model, and the superscript of $\bm{u},\hat{\bm{u}}$ denotes their time steps in the segment.
@@ 4 Experiments
Here, we present comprehensive evaluation of our methods. Firstly, we validate the feasibility and efficacy of the Laplace block to compute the Laplacian on coarse non-uniform meshes. Subsequently, the entire model is trained and evaluated on various PDE datasets and compared with several baseline models. Finally, we perform ablation studies to demonstrate the positive contribution of each individual component within our model and the impact of numerical integrator. The source code is publicly available in the Pytorch [^1] https://github.com/intell-sci-comput/PhyMPGN and Mindspore [^2] https://gitee.com/mindspore/mindscience/tree/master/MindFlow/applications/data_mechanism_fusion/phympgn repositories.
@@ 4.1 Laplace block validation
We create a synthetic dataset to validate the effectiveness of the Laplace block. First, we generate 10,000 samples of random continuous functions on a high-resolution ($129\times 129$) regular grid by cumulatively adding sine and cosine functions, satisfying periodic BCs. By employing the 5-point Laplacian operator of the finite difference method, the Laplacian of the functions at each node is computed as the ground truth. Then the dataset is down-sampled from the high-resolution regular grids to a set of coarse non-uniform mesh points yielding only 983 nodes. Note that 70% of the dataset is used as the training set, 20% as the validation set, and 10% as the test set. More details of the synthetic dataset are shown in Appendix D.1.
We compare the performances of four methods: (1) Laplace block with BC padding; (2) the discrete Laplace-Beltrami operator with BC padding, referred to as Mesh Laplace; (3) the spatial difference layer for Laplacian, referred to as SDL (Seo et al., 2020); (4) SDL with BC padding, referred to as SDL-padding. The MSE and RNE of all experiments are shown in Table 2. Mesh Laplace performs poorly compared to other learnable methods, while our Laplace block with the fewest parameters greatly outperforms other methods. SDL-padding exhibits an obvious performance improvement, indicating the effectiveness and importance of our padding strategy. A snapshot of a random function, ground truth of its Laplacian, and prediction are shown in Appendix Figure A.4.
Table 1: The MSE and RNE of four methods to approximate Laplacian on coarse domain.
@@ 4.2 Modeling PDE systems
Solving PDEs serves as a cornerstone for modeling complex dynamical systems. However, in fields such as climate forecasting, reactions of new chemical matters, and social networks, the underlying governing PDEs are either completely unknown or only partially known. Hence, there is a great need for data-driven modeling and simulation. Based on the assumption, we train and evaluate our model on various physical systems, including Burgers’ equation, FitzHugh-Nagumo (FN) equation, Gray-Scott (GS) equation, and flows past a cylinder, particularly in small training data regimes. We also compare our model against several representative baseline models, including DeepONet (Lu et al., 2021), PA-DGN (Seo et al., 2020), MGN (Pfaff et al., 2021), and MP-PDE (Brandstetter et al., 2022). The numbers of learnable parameters and training iterations for all models are kept as consistent as possible to ensure a fair comparison. All the data are generated using COMSOL with fine meshes and small time stepping, which are further spatiotemporally down-sampled to establish the training and testing datasets. More details of the datasets are discussed in the Appendix D.2.
@@
Figure 5: Four scenarios with different domains and hybrid BCs.
@@
To validate our model’s ability to encode different types of BCs, we use Burgers’ equation as an example and consider three additional scenarios with irregular domain and hybrid BCs, resulting in total four scenarios, as illustrated in Figure 5: (1) Scenario A, a square domain with periodic BCs, as mentioned in the previous paragraph; (2) Scenario B, a square domain with a circular cutout, where periodic and Dirichlet BCs are applied; (3) Scenario C, a square domain with a dolphin-shaped cutout, where periodic and Dirichlet BCs are applied; and (4) Scenario D, a square domain with semicircular cutouts on both the left and right sides, where periodic and Neumann BCs are applied. Similarly, 10/10 training and testing trajectories are generated and down-sampled with nearly $N=600$ observed nodes in each domain. Figure 6 depicts the error distribution over all time steps, correlation curves, and predicted snapshots at the last time step, for the last three scenarios (results of Scenario A are shown in Figure 4a). It is seen that our model significantly outperforms baseline models and shows the efficacy of handling PDE systems with different BCs.
@@
Computational efficiency:
The computational cost of the Laplace block and the padding strategy is minimal. The computation of Mesh Laplace in the Laplace block includes only matrix multiplication, while the cost of the padding strategy involves merely copying the corresponding nodes in the domain and is small. A detailed analysis and comparison with MGN and MP-PDE as well as the classical numerical method (COMSOL) are presented in Appendix E. In all experiments, our model outperforms the baselines with considerable margins, e.g., exceeding 50% gains (see the specific numerical results reported in Appendix Tables A.5, A.7 and A.8). The predicted trajectory snapshots of our model and ground truth are provided in Appendix Figure A.6.
Figure 8: Results of Models A, B, C, and Ours.